### v1

# You Only Acquire Sparse-channel (YOAS): A Unified Framework for Dense-channel EEG Generation

Abstract
High-precision acquisition of dense-channel electroencephalogram (EEG) signals is often impeded by the costliness and lack of portability of equipment. In contrast, generating dense-channel EEG signals effectively from sparse channels shows promise and economic viability. However, sparse-channel EEG poses challenges such as reduced spatial resolution, information loss, signal mixing, and heightened susceptibility to noise and interference. To address these challenges, we first theoretically formulate the dense-channel EEG generation problem as by optimizing a set of cross-channel EEG signal generation problems. Then, we propose the YOAS framework for generating dense-channel data from sparse-channel EEG signals. The YOAS totally consists of four sequential stages: Data Preparation, Data Preprocessing, Biased-EEG Generation, and Synthetic EEG Generation. Data Preparation and Preprocessing carefully consider the distribution of EEG electrodes and low signal-to-noise ratio problem of EEG signals. Biased-EEG Generation includes sub-modules of BiasEEGGanFormer and BiasEEGDiffFormer, which facilitate long-term feature extraction with attention and generate signals by combining electrode position alignment with diffusion model, respectively. Synthetic EEG Generation synthesizes the final signals, employing a deduction paradigm for multi-channel EEG generation. Extensive experiments confirmed YOAS’s feasibility, efficiency, and theoretical validity, even remarkably enhancing data discernibility. This breakthrough in dense-channel EEG signal generation from sparse-channel data opens new avenues for exploration in EEG signal processing and application.
Index Terms:
Multi-channel, Sparse-channel, Dense-channel, EEG signal generation, Transformer, Generative adversarial network, Diffusion model, Neuroscience, Brain-computer interface
I Introduction
Electroencephalogram (EEG), crucial for neuroscience and clinical diagnosis, captures neuronal electrical activity from the cerebral cortex. In research, EEG offers new insights into brain function, elucidating cognitive processes, sleep patterns, neural mechanisms of brain disorders, etc [1, 2, 3, 4]. In clinical settings, EEG aids in diagnosing epilepsy [5], sleep disorders [6], and informs surgical decisions [7]. It also can assesse treatment efficacy [8] and contribute to advancements in brain-computer interface (BCI) technology [9], enhancing rehabilitation and quality of life for individuals with disabilities.
High-precision dense-channel EEG acquisition is essential across diverse applications, yet the cost and portability limitations of dense-channel EEG devices pose significant challenges. A raised question is that is it feasible to develop a framework capable of generating multi-channel EEG signals from a limited number of channels?
Neural activity is influenced by factors such as neuron types, synaptic connections, and neural circuits [10]. And as a technique for exploring neural activity, EEG are high-dimensional and dynamic, necessitating the capture of intricate temporal correlations and spatial distributions [11, 12]. Therefore, generating EEG signals presents challenges in simulating the complex brain activity, which is characterized by the spatiotemporal properties of each channel. Moreover, generated EEG signals exhibit uncertainty and variability due to physiological and environmental factors, posing limitations to model training and performance evaluation [13, 14]. In summary, the generation of multi-channel EEG signals faces the following challenges:
Firstly, as each electrode corresponds to specific brain areas, harmonizing signal characteristics from electrodes at different locations under the same mental condition presents challenges. This is due to that the distinct characteristics of EEG signals vary among individuals across different inner states [15], and the electrode signals exhibit unique features at various positions [16, 17].
Secondly, EEG signal generation for whole brain only based on few channels is exceptionally challenging. Initially, the low spatial resolution of EEG signals makes pinpointing their sources quite difficult. Further, within the International 10-20 system, the electrode signals in the neighboring areas may have weak correlation, while those from distant areas may exhibit strong correlation. Additionally, the previous studies [18] have only focused on generating EEG data tied to specific stimuli.
Thirdly, the generation models of EEG signals pose the challenges including training instability, inefficiency, and susceptibility to signal noise. Despite the advancement of generation models such as Variational Auto Encoder (VAE) [19] and Generative Adversarial Network (GAN) series [20] in general domain, these issues persist for the specific EEG signal generation problem.
Figure 1: The YOAS framework within the exploratory scheme includes Data Preparation, extracting Bias Signal from Raw Data through regional division, followed by Data Preprocessing to eliminate outliers. The Biased-EEG Generation module captures long-term EEG features and aligns spatial positions to generate Two-stage Bias. Finally, the Synthetic EEG Generation module synthesizes the final EEG signal.
Therefore, to address the aforementioned challenges, we proposed a specialized model for dense-channel EEG generation problem, named YOAS comprising four sequential modules, depicted in Fig. 1, to improve the efficiency of EEG signal generation for whole brain:
Data Preparation: Following the International 10–20 system, EEG electrode distribution reflects approximately hemispherical symmetry and functional segregation of human brain [21, 22]. Electrodes are initially grouped into distinct electrode regions, then the reference channel signal within each region is selected, and finally is subtracted from other channels within the same region to derive the EEG bias signal for each electrode.
Data Preprocessing: In the context of EEG bias signals derived from the Data Preparation stage, we commence by empirically eliminating data points within each signal segment deviating by one or two standard deviations from the total segment signal of each channel. Subsequently, linear interpolation is employed to address the missing data points. Finally, Multiscale Principal Components Analysis (MSPCA) [23] is utilized to reduce dimensionality and extract features from the interpolated data.
Biased-EEG Generation: The Biased-EEG Generation modules comprise BiasEEGGanFormer and BiasEEGDiffFormer. BiasEEGGanFormer captures high timing resolution characteristics using Transformer [24] with multi-head attention, producing the “One-stage Generated Bias”. In BiasEEGDiffFormer, spatial alignment and feature extraction are performed using ConvBlock. DiffusionBlock then generates signal samples incrementally from noise data, ensuring consistency with electrode locations, resulting in the “Two-stage Generated Bias”.
Synthetic EEG Generation: In this module, we synchronize the Regional Reference Signal with the Two-stage Generated Bias to achieve spatial alignment. Then, we add the Regional Reference Signal to the Two-stage Generated Bias to generate signals for each channel. To cover the entire brain, we adopt divide-and-conquer approach, selecting key channels within each region as references and optimizing the derivation path with greedy algorithm.
In summary, this study contributes significantly in the following two aspects:
Formulation of the Dense-channel EEG Generation Problem: We formulate the problem of generating dense-channel EEG signals from sparse channels as a cross-channel EEG generation task. This formulation addresses challenges related to biased EEG signal generation, EEG electrode position alignment and the synthesis of EEG signals across channels.
Development of a Unified Framework for Dense-channel EEG Generation: We introduce YOAS, a unified framework for dense-channel EEG generation. This framework integrates prior information on electrode positioning from the International 10-20 system, precise bias EEG data for each channel, and general across-channel EEG synthesis paradigms to effectively produce multi-channel EEG signals.
II RELATED WORK
II-A Across-channel EEG Signal Generation
In this study, we delineate the cross-channel EEG signal generation task as the process of generating an additional EEG channel signal based on the collected EEG channel signal. Firstly, previous studies predominantly concentrated on data augmentation, but failed to accomplish cross-channel EEG signal generation. For instance, EEG-GAN [20] primarily targets the single-channel data augmentation, showing an enhancement over WGAN [25]; Nonetheless, its pursuit of capturing features in lengthy sequences and local intricacies escalates computational costs and training instability significantly. Moreover, although BioDiffusion [26] excels in producing high-fidelity multivariate signals, it does not specifically cater to cross-channel EEG signal generation. Consequently, we introduce the Biased-EEG Generation module incorporating Transformer [24] and diffusion model [27, 28], thereby facilitating cross-channel EEG signal generation.
II-B EEG Electrode Position Alignment
The EEG electrode position alignment refers to the process of aligning generated EEG signal data with the physical positions of the electrodes to which they correspond. Generally speaking, diffusion models [27] operating by incrementally introducing and reducing noise through both forward and reverse processes, is widely used to generate the data in specific conditions. The introduction of DDPM [28] further enhances the model, surpassing GAN [29] in generating high-quality images. While diffusion models have been applied to generate high-fidelity, non-stationary biomedical signals [26], they primarily focus on enhancing single-channel data, not multi-channel or dense-channel data. Thus, despite the advancements, a crucial issue that remains unresolved is the lack of adherence to the positional characteristics of the electrodes in the generated signals. In BiasEEGDiffFormer of YOAS, we propose an innovative diffusion model to align electrode positions by matching signals generated by BiasEEGGanFormer with corresponding reference signals in their respectively divided regions.
II-C Across-channel EEG Synthetic Paradigm
The across-channel EEG synthetic paradigm refers to the deduction paradigm process that pursuing the optimized sets of the generation paths for generating multi-channel EEG signals across the whole brain. As it stated before, most of EEG signal generation studies primarily focused on signal generation for data augmentation [10, 26]. However, in the pursuit of generating multi-channel EEG signals throughout the entire brain, the task of deriving EEG signals for all channels from a single channel signal frequently presents significant challenges. Hence, it becomes imperative to determine and refine the optimal number of reference channel signals required to generate signals across all channels. Additionally, the optimization and delineation of generation pathways for specific channel signals based on the designated reference channel signals are also paramount. In response to these challenges, we introduce the Synthetic EEG Generation module within the YOAS framework to effectively address the aforementioned concerns.
III METHOD
III-A Overview
The task for dense-channel EEG signal generation for entire brain using YOAS, as depicted in Fig. 1, could be theoretically formulated by optimizing a set of across-channel EEG signal generation problems based on $K$ reference channels, where each reference channel represents a regional division (RD). Initially, as shown in Data Preparation stage of Fig. 1, the entire brain positions are partitioned into $K$ RDs according to the widely-used International 10-20 system in EEG, denoted as $RD_{i}$ for the $i$-th division and $M_{i}$ as the total number of channels within $RD_{i}$. Secondly, we can record the channel number of the $i$-th RD as the set $\hat{M}_{i}$, $\hat{M}_{i}=\{1,2,\ldots,j,\ldots,M_{i}\}$, with $j$ standing for the $j$-th channel ($1<j\leq M_{i}$) in the $RD_{i}$. Thirdly, the whole brain channel signal set can be formulated as $\{O_{ij}\}_{i=1,j=1}^{K,M_{i}}$, the reference channel signal set can be defined as $\{O_{i1}\}_{i=1}^{K}$ by selecting the first channel for each RD and the remaining channel signal set can be represented by $\{O_{ij}\}_{i=1,j=2}^{K,M_{i}}$. Thus, the across-channel EEG signal generation problem is involved by asking that “can the remaining channel signals $O_{ij}$ ($1<j\leq M_{i}$) in the $RD_{i}$ be accurately generated based on the reference channel $O_{i1}$, subject to $0<i\leq K$?” Mathematically, the dense-channel EEG signal generation problem of whole brain can be initially expressed as follows:
$Gen_{n}(O_{i1})\to C_{ij},\$ (1)
$\{Gen\}=\{Gen_{1},Gen_{2},...,Gen_{n},...,Gen_{N}\},$ (2)
where $\{Gen\}$ represents a set of $N$ possible strategy to generate signals, and $C_{ij}$ represents the generated signal of $O_{ij}$, satisfying $i\in[1,K]\;,\;j\in[2,M_{i}]\;,n\in[1,N]$.
In order to accurately solve Formula (1) and Formula (2), here we innovatively introduce the variable of the bias signal of each regional division, namely, $RD_{i}$, based on the corresponding reference signal $O_{i1}$ ($0<i\leq K$), which can be calculated as follow:
$\{\beta_{ij}\}_{i=1,j=2}^{K,M_{i}}=\{O_{ij}\}_{i=1,j=2}^{K,M_{i}}-\{O_{i1}\}_{i=1}^{K},$ (3)
where $\beta_{ij}$ represents the bias data of other channel signals in $RD_{i}$ relative to the reference channel. Subsequently, we could rewrite Formula (1) and Formula (2) respectively as:
$Gen^{\prime}_{n}(O_{i1},\beta_{ij})\to C_{ij},\$ (4)
$\{Gen^{\prime}\}=\{Gen^{\prime}_{1},Gen^{\prime}_{2},...Gen^{\prime}_{n}...,Gen^{\prime}_{N}\}.$ (5)
To address the optimization of Formulas (4) and (5), the key issue lies on refining the generation process of bias signals. In this study, the quality of the bias signal entails Data Preprocessing stage as shown in Fig. 1 for $\beta_{ij}$, elaborated in Section III-C, resulting in $\beta^{\prime}_{ij}$.
Subsequently, as shown in Biased-EEG Generation stage of Fig. 1, we incorporate the temporal attributes of the biased EEG signal and the spatial characteristics of the electrode positions, organizing the generation process into two stages. This entails the utilization of two generation modules, labeled as $Trans$ (illustrated in Fig. 3) and $Diff$ (depicted in Fig. 4), to produce One-stage bias $b_{ij}$ and Two-stage bias $B_{ij}$, respectively. Thus, the procedure is expressed as follows, with details provided in Section III-D:
$b_{ij}$ $=Trans(\beta^{\prime}_{ij}),$ (6)
$B_{ij}$ $=Diff(b_{ij}).$ (7)
Finally, as displayed in Synthetic EEG Generation stage of Fig. 1, we synthesize the Two-stage bias $B_{ij}$ and the reference channel $O_{1j}$ of the $RD_{i}$ to generate $C_{ij}$, satisfying $0<i\leq K$ and $0<j\leq M_{i}$. Taking into account Formula (4), we further have expressions as follows:
$C_{ij}=yoas(B_{ij},O_{i1},Gen^{\prime}_{n}(O_{i1},B_{ij})),$ (8)
where the $yoas(.)$ means a final EEG generation operation that $B_{ij}$ is combined with $O_{1j}$ according to the generation path strategy of $Gen^{\prime}_{n}(O_{i1},B_{ij})$ to generate $C_{ij}$.
Once the channel signal in $RD_{i}$ has been generated, the corresponding generation path within this division is simultaneously determined. However, the feasibility of merging the corresponding generation paths across regional divisions remains uncertain. Therefore, we execute the merging process between the generated paths and rewrite Formula (8) accordingly as follows:
$X=bool(Gen^{\prime}_{n}(O_{m1},\beta_{i1})\to C_{i1}),$ (9)
$C_{ij}=yoas^{\prime}(B_{ij},\{O_{i1},O_{m1}\},X,Gen^{\prime}_{n}(O_{i1},B_{ij}),$ (10)
$C_{ij}=\hat{yoas}(B_{ij},\{O_{i1},O_{m1}\}$ (11)
$Gen^{\prime}_{n}(O_{i1},B_{ij}),Gen^{\prime}_{\hat{n}}(O_{m1},B_{ij})),$
$\hat{RD}_{t}=RD_{i}\oplus RD_{m}.$ (12)
[…]
III-B Data Preparation
This section elucidates the initial regional division discussed in Section III-A, which divides all brain channels into $K$ RDs using the International 10–20 system of EEG electrode arrangement and the approximately structural/functional symmetry between the left and right hemispheres [16]. To achieve this, we propose three hypotheses for multi-channel EEG generation pathways, illustrated in Fig. 2. For convenience, we denote that $O_{ij}$ represents the source channel, and $O_{im}$ represents the target channel within the $RD_{i}$. Moreover, $Dis(.)$ calculates the physical distance between two electrodes, while $D(.)$ computes the correlation distance between their signals, where $D=\mid 1-\rho\mid$ $(-1<\rho<1)$, with $\rho$ representing the Pearson Correlation Coefficient.
Hypothesis 1. In Fig. 2(A), the signal of a given channel may be generated by another nearby channel. If the criteria
$\begin{cases}Dis(O_{ij},O_{im})\leq L_{1}\\ D(O_{ij},C_{im})\leq P_{1}\end{cases}$ is met, it implies that the signal of the target channel $O_{im}$ can be generated through the source channel $O_{ij}$, i.e., $O_{ij}\longrightarrow O_{im}$.
Hypothesis 2. In Fig. 2(B), for a specific channel, its signal may be indirectly derived from another channel within the same $RD_{i}$ through intermediate channel $O_{in}$. If the condition
$\begin{cases}\begin{cases}Dis(O_{ij},O_{in})\leq L_{2}\&Dis(O_{ij},O_{im})\leq L_{2}\\ \&Dis(O_{im},O_{in})\leq L_{2}\end{cases}\\ D(O_{ij},C_{in})\leq P_{2}\\ D(O_{in},C_{im})\leq P_{2}\end{cases}$ is satisfied, it indicates that the signal of target channel $O_{im}$ can be indirectly generated by that of the source channel ${O_{ij}}$, i.e., $O_{ij}\xrightarrow{o_{in}}O_{im}$.
Hypothesis 3. In Fig. 2(C), for channels with certain level of symmetry between the left and right hemispheres of the brain, their signals can mutually generate. Thus, if the criteria $\begin{cases}Dis(O_{ij},O_{im})\leq L_{3}\\ D(O_{ij},C_{im})\leq P_{3}\\ D(C_{ij},O_{im})\leq P_{3}\end{cases}$ is satisfied, this signifies a mutual signal generation between the source channel $O_{ij}$ and the target channel $O_{im}$, i.e., $O_{ij}\longleftrightarrow O_{im}$.
To this end, we define the maximum physical radiation generation range set of the reference channel in the above three hypotheses as $L=\{L_{1},L_{2},L_{3}\}$, and the coefficient set of the maximum feasible generation signal of the reference channel signal is recorded as $P=\{P_{1},P_{2},P_{3}\}$, $\left.\left(\begin{matrix}0\leq P_{i}\leq 2\\ \end{matrix}\right.\right)$. In the International 10-20 system for EEG placement, zoning proceeds as follows: initially, we empirically partition it into 5 regions comprising the Frontal Lobe, Temporal Lobe, Central Sulcus, Parietal Lobe, and Occipital Lobe. Subsequently, leveraging the approximately hemispherical symmetry, we utilize the central sulcus as the axis for division, excluding paired regions with notable signal correlation such as (A1, A2) and (O1, O2). Further, we apply the aforementioned hypotheses to refine the regional division task. Hypothesis 1 (III-B) sets $L_{1}$ empirically to the radius of the 2D distribution map with $P_{1}$ equal to 0.3. Hypothesis 2 (III-B) maintains $L_{2}$ at the 2D distribution map’s radius, with $P_{2}$ equal to 0.3. Hypothesis 3 (III-B) sets $L_{3}$ to the 2D distribution map’s diameter, with $P_{3}$ equal to 0.1.
III-C Data Preprocessing
Given the low signal-to-noise ratio (SNR) of EEG signals and the potential for signal corruption, it is likely that the bias signal data $\{\beta_{ij}\}_{i=1,j=2}^{K,\hat{M}}$ derived from the Data Preparation stage of Fig. 1 may contain erroneous data. To mitigate these challenges, the following preprocessing procedures have been implemented. Initially, outlier removal is conducted by empirically eliminating data points within each signal segment that deviate by one or two standard deviations from the total segment signal of each channel. Subsequently, linear interpolation is applied to address missing data points. Additionally, NaN values in $\{\beta_{ij}\}_{i=1,j=2}^{K,\hat{M}}$ are filled using linear interpolation, and any remaining INF values post-interpolation are replaced with zero. Finally, Multiscale Principal Components Analysis (MSPCA) [23] is employed to reduce dimensionality and extract effective features from the interpolated data.
III-D Biased-EEG Generation
Previous methodologies predominantly focused on signal generation for single-channel data augmentation, thereby exhibiting limitations in the context of dense-channel EEG signal generation. This task presents challenges in terms of efficiency and consideration of electrode position characteristics. To overcome these constraints, we introduce the Biased-EEG Generation module, consisting of two sub-modules: BiasEEGGanFormer (shown in Fig. 3) and BiasEEGDiffFormer (displayed in Fig. 4).
Figure 2: Illustration of hypotheses for multi-channel EEG generation pathways: (A) Direct one-way generation involves generating signals directly with reference to the channel; (B) Indirect one-way generation requires an intermediate channel for the reference channel to generate signals indirectly; (C) Mutual generation involves two channels serving as reference channels for each other, facilitating mutual signal generation.
Figure 3: BiasEEGGanFormer employs an adversarial generation architecture, where the Biased-EEGGenerator generates data fed into the Biased-EEGDiscriminator for loss feedback, thereby improving the generation process quality. The resulting output comprises the One-stage Generated Bias, enhancing EEG signal refinement.
Figure 4: BiasEEGDiffFormer aligns One-stage Generated Bias with the correspondingly physical locations, first applying one-dimensional convolution in ConvBlock for feature extraction, and subsequently involving DiffusionBlock to generate Two-stage Generated Bias based on Regional Reference Signal and a diffusion model.
Figure 5: The 32-channel EEG 2D distribution map and the initial EEG channel regional division, with Fp1, Fz, Cz, A1, CP1, CP2, Pz, T5, Oz as reference channels for their respectively regional division.
1) BiasEEGGanFormer
Compared with the WGAN series [30], the Transformer model [24] adopts a self-attention mechanism, enabling it to capture long-term dependencies in time series data without being constrained by sequence length. Meanwhile, compared with traditional GAN models, the Transformer architecture mitigates issues such as generator collapse and training instability [30]. To generate quasi-realistic data, we propose BiasEEGGanFormer as shown in Fig. 3, which retains the original GAN architecture while incorporating the Transformer. Specifically, inspired by [31], we opt for a simpler approach, adding linear and ReLU layers after the Transformer structure. For the generator, an embedding step extracts features from input data, followed by the Multi-Head Self-Attention module capturing sequence dependencies. Subsequently, an MLP with two linear layers and ReLU activation further processes encoded and decoded outputs, with data generated through a fully connected layer. Both generator and discriminator share the same structure for balance, producing data termed as ’One-stage Generated Bias’ i.e. $b_{ij}$ described in Formula 6. Unlike previous works [30],[32],[33] employing numerous convolutional layers or complex up-sampling and down-sampling techniques, our BiasEEGGanFormer prioritizes the generation effectiveness and efficiency.
2) BiasEEGDiffFormer
Taking advantage of the correlation between EEG signals, electrode positions, and reference channel signals, BiasEEGDiffFormer (Fig. 4) is developed as an extension of the BiasEEGGanFormer. Initially, we align each One-stage Generated Bias $b_{ij}$ with the reference signal $O_{i1}$ of its associated physical location $RD_{i}$. By merging these generated data with reference signals, we construct a virtual two-dimensional EEG distribution map, which enhances the interpretability of generated data $C_{ij}$ for each channel. Next, we utilize the ConvBlock to further extract features, followed by DiffusionBlock [28]. Drawing upon the regional reference signal, we meticulously select the generated data that most accurately aligned with its position, resulting in ’Two-stage Generated Bias’ i.e. $B_{ij}$ describled in Formula 7 with a more detailed expression as
$B_{ij}=Diff(b_{ij},O_{i1},RD_{ij}).$ (13)
Typically, the forward process of the diffusion model gradually adds noise to the original signal to generate a noise signal. For each time step $t$, the change in the signal can be expressed as:
$B_{i,j,t}$ $=\sqrt{\bar{\alpha}_{t}}\cdot B_{i,j,0}+\sqrt{1-\bar{\alpha}_{t}}\cdot\epsilon,$ (14)
$q(B_{t}|B_{i,j,t-1})$ $=\mathcal{N}(B_{i,j,t};\sqrt{\alpha_{t}}B_{i,j,t-1},(1-\alpha_{t})I),$
$B_{i,j,0}$ $=Diff_{init}(b_{ij},O_{i1},RD_{ij}),$
$\bar{\alpha}_{t}$ $=\prod_{s=1}^{t}\alpha_{s},$
$\epsilon$ $\sim\mathcal{N}(0,I),$
where $B_{i,j,0}$ is the initial signal, $\bar{\alpha}_{t}$ is the cumulative diffusion factor of time step $t$, and $\epsilon$ is the noise of standard normal distribution. Thus, the formulation of the forward diffusion process can be expressed as:
$q(B_{i,j,t}|B_{i,j,0})$ $=\mathcal{N}(B_{i,j,t};\sqrt{\bar{\alpha}_{t}}\cdot Diff_{\mathrm{init}}(b_{ij},O_{i1},RD_{i}),$ (15)
$(1-\bar{\alpha}_{t})I).$
The backward process is used to gradually recover the original signal from the noisy signal. For each time step $t$, the backward process can be expressed as:
$p_{\theta}(B_{i,j,t-1}|B_{i,j,t})=$ $\mathcal{N}(B_{i,j,t-1};\mu_{\theta}(B_{i,j,t},t),\sigma_{t}^{2}I),$ (16)
$\mu_{\theta}(B_{i,j,t},B_{i,j,t})=$ $\frac{1}{\sqrt{\alpha_{t}}}\left(B_{i,j,t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(B_{i,j,t},t)\right),$
where $\mu_{\theta}(B_{i,j,t},t)$ is used to predict the mean, and $\sigma_{t}^{2}$ is the variance of time step $t$. To train the neural network $\epsilon_{\theta}$, the mean square error is used as the loss function:
$L=\mathbb{E}_{t,B_{i,j,0},\epsilon}\left[\|\epsilon-\epsilon_{\theta}(B_{i,j,t},t)\|^{2}\right],$ (17)
by repeatedly applying the backward denoising process, where the final signal $B_{ij}$ is gradually generated from the noise. Thus, the generation process of $B_{i,j}$ can be rewritten to Algorithm 1 as below.
Algorithm 1 Generation Process of $B_{ij}$
0:  $b_{ij}$, $O_{i1}$, $RD_{i}$, $P_{i}$, $\hat{t}$, $T$
0:  Generated signal $B_{ij}$
  Initialization: $B_{i,j,0}=Diff_{\text{init}}(b_{ij},O_{i1},RD_{i})$
  Forward Process: Generate noisy signal $B_{i,j,t}$ for $t=1,\ldots,T$:
$B_{i,j,t}=\sqrt{\bar{\alpha}_{t}}\cdot B_{i,j,0}+\sqrt{1-\bar{\alpha}_{t}}\cdot\epsilon$
where $\epsilon\sim\mathcal{N}(0,I)$.
  Backward Process:Iteratively denoise from $t=T$ to $t=1$:
$\small B_{i,j,t-1}$ $\sim\mathcal{N}(\mu_{\theta}(B_{i,j,t},t),\sigma_{t}^{2}I)$
$\mu_{\theta}(B_{i,j,t},t)$ $=\frac{1}{\sqrt{\alpha_{t}}}\left(B_{i,j,t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(B_{i,j,t},t)\right)$
$\sigma_{t}^{2}$ $=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}(1-\alpha_{t})$
$\epsilon_{\theta}(B_{i,j,t},t)$ $\approx\text{NN}(B_{i,j,t},b_{ij},O_{i1},RD_{i},t)$
$\textbf{If}\;D_{t}(B_{i,j,t},$ $O_{i,1},RD_{i})\leq P_{i}\;\&\;\underset{i\in T}{\textbf{argmin}}\;D_{i}=D_{t}$
$\textbf{then}\;\hat{t}$ $=t$
  Loss Function: Compute the loss for training:    $L=\mathbb{E}_{t,x_{0},\epsilon}\left[\|\epsilon-\epsilon_{\theta}(B_{i,j,t},t)\|^{2}\right]$
  Return: $B_{i,j,\hat{t}}$
Algorithm 2 : Synthetic Paradigm 1
[…]
0:  Refined generation paths among channels //Direct,Indirect,Mutual
  start
  for $Ch_{i}$ in Channel list do
     for $Ch_{j}$ in Channel list $(i\neq j)$ ($j\in Z^{*}\&1\leq j\leq N$do
        Set $Ch_{i}$ as Reference Channel
        if $Ch_{i}\to Ch_{j}\&Ch_{i}\to Ch_{j+n}$ $n\in([1-j,-1]\cup[1,N-j]),n\in Z$ then
           $Ch_{i}\to Ch_{j}\mid Ch_{j+n}$   //Hypothesis 1(III-B)
        else if $Ch_{i}\to Ch_{j}\&Ch_{j}\to Ch_{m}\;\;(i\neq j\neq m)$ then
           $Ch_{i}\to Ch_{m}$      //Hypothesis 2(III-B)
        else if $Ch_{i}\to Ch_{j}\&Ch_{j}\to Ch_{i}$ then
           $Ch_{i}\leftrightarrow Ch_{j}$       //Hypothesis 3(III-B)
        else
           return  ”Not feasible”
        end if
     end for
  end for
Algorithm 3 : Synthetic Paradigm 2
[…]
0:  Merge results
  start
  if Reference Channel of Regional Division A is $Ch_{i}$ Reference Channel of Regional Division B is $Ch_{j}$ $Ch_{i}\to Ch_{j}\;\&\;Ch_{j}\to Ch_{i}$ then
     Merge Regional Division A and Regional Division B
  else
     return  ”No Merge”
  end if
TABLE I: Parameters setting for the training/testing stage of YOAS.
Parameter Value
Input size 7500
Output size 7500
Hidden size 512
Optimize Adam
Learning rate 0.0001
Number of epochs with early stopping 10000
Batch size 1024
Learning rate decay 0.95
Num layers 6
Num heads 8
III-E Synthetic EEG Generation
As shown in Fig. 1, to pursue optimized sets of generation paths for generating multi-channel EEG signals across the whole brain, we propose a deduction paradigm consisting of two consecutive paradigms: i.e, Synthetic Paradigm 1 (referring to Algorithm 2) and Synthetic Paradigm 2 (referring to Algorithm 3). Given the initial regional division’s insufficient accuracy in Section III-B Data Preparation, Synthetic Paradigm 1 aims to refine signal generation paths, including direct, indirect, and mutual generation among channels. Building upon the refined results of Synthetic Paradigm 1, Synthetic Paradigm 2 is designed to retrieve the final optimized sets of the generation paths and regional divisions for generating dense-channel EEG signals across the whole brain.
IV EXPERIMENTS
IV-A Implementation Details
Datasets: We conduct training and validation of our models on the FACED dataset provided in reference [34]. The dataset comprises recordings from 123 subjects, each with 32 EEG channels, while they watched 28 video clips, all lasting for 30 seconds. These clips cover a range of emotions, where we select four clips for neutral emotion. In the experiments, we select 250Hz as the sampling frequency. There are a total of 55 subjects, including 30 subjects for training, and 25 subjects for testing.
Training Details: We conduct training and testing on the provided dataset using the same parameters outlined in Table LABEL:tab:table1. Initially, for the data utilized during training, we employ WGAN-GP [30] to assess the feasibility of each generation path showed in Fig. 5 through the testing of one-to-one channel signal generation. Subsequently, compared with WGAN-GP, we retest the performance of the paths generated by our BiasEEGGanFormer. All computations are executed on NVIDIA Tesla V100, and the Early-Stopping strategy is applied in the training stage.
IV-B Performance Comparison
[…]
Figure 6: Performance comparison between BiasEEGGanFormer (orange) and WGAN-GP (blue): (A) and (B) demonstrate that BiasEEGGanFormer takes only 22.22% of the time cost of WGAN-GP, while maintaining comparable generation quality. (C) and (D) provide further evidence, showing that BiasEEGGanFormer outperforms WGAN-GP in quality across 41 generation path experiments, with less time cost.
Figure 7: The power spectrum comparison between the “indifferent generated signal” of WGAN-GP in (B), and BiasEEGGanformer in (C), demonstrates BiasEEGGanformer’s superior performance in attenuating high-amplitude components in frequency-spectrum domain, potentially indicative of noise.
Figure 8: Multi-channel EEG signal generation paradigm on neutral emotion data. (A), (C), and (D) are the generation paradigms with Fp1, Fz, and Pz as reference channels, respectively. (B) shows that CP1 and CP2 as reference channels in one-way generation paradigms alone, while pair of O1 and O2 and pair of A1 and A2 are the mutual generation paradigms.
IV-C Optimized Generation Paths
Based on the final output of the YOAS framework in our experiments, we observed that the initial ten regional divisions depicted in Fig. 5 were optimized and merged into seven regional divisions on neutral emotion data, as demonstrated in Fig. 8 following Synthetic Paradigm 2. Specifically, as illustrated in Fig. 8(A), Fp1 can directly generate F3, F7, FC5, and Fp2, while FP2 indirectly generates F4, F8, and FC6 through Fp1. Notably, F7 can directly generate FC5, and F8 can directly generate FC6. In Fig. 8(B), CP1 generates T3 exclusively, CP2 generates T4 exclusively, while pair of O1 and O2 reciprocally generate each other, as well as pair of A1 and A2, can generate each other mutually. Referring to Fig. 8(C), Fz directly generates FC1, FC2, and Cz, while indirectly generating C3 and C4 through Cz. As shown in Fig. 8(D), Pz directly generates Cp5, Cp6, P4, and PO4, with indirect generation of P3 through P4. Notably, PO4, PO3, and Oz mutually generate each other, with Oz also generating T5 and T6, both of which mutually generate each other.
V discussion and Conclusion
We firstly theoretically formulate the problem of generating dense-channel EEG signals from sparse channels by optimizing a set of cross-channel EEG generation tasks. We further introduce the novel unified framework YOAS for dense-channel EEG generation, leveraging prior information on electrode positioning from the International 10-20 system EEG placement, precise bias EEG data for each channel, and general across-channel EEG synthesis paradigms. Additionally, we validate three conjectures of generation pathways (depicted in Fig. 2), contributing to the advancement of dense-channel EEG signal generation methodologies and pioneering new frontiers in this field.
The proposed YOAS currently falls within the realm of data-driven statistical models, where parameters $L$ and $P$ in Data preparation stage (III-B) of YOAS are empirically set, and their correlation with physical mechanisms remains to be investigated in future research. Also, the threshold values for outlier detection are predominantly determined empirically, thus, future work may explore the adaptive detection of outliers in Data Preprocessing stage (III-C) of YOAS, drawing inspiration from metrics such as the Interquartile Range (IQR) [35].
Our innovative approach revolutionizes the construction of dense-channel EEG systems, promising broader applications. By optimizing EEG systems to support more channels with fewer resources, we potentially significantly reduce the cost of high-density electrode arrays while maintaining its performance. This breakthrough enhances EEG accessibility in healthcare, consumer electronics, and neuroscience.